OpenStudy (anonymous):

4*(a/b)=a/(4b) True or false

OpenStudy (anonymous):

I need this answered so I can use it as an example for other questions

OpenStudy (mertsj):

False

OpenStudy (mertsj):

\[4(\frac{a}{b})=\frac{4}{1}(\frac{a}{b})=\frac{4a}{b}\]